OpenStudy (anonymous):

The distance from Earth to the Moon is approximately 240,000 mi. The speed of light is 3.00 * 10^8 m/s. How long does it take for light to travel from Earth to the Moon and back again? Please Show work.

OpenStudy (anonymous):

time from the earth to the moon = 240000m /(3.00*10^8)m/s time = .0008s therefore to the moon and back = .0016s

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

well its miles to meters

OpenStudy (anonymous):

so first have to convert

OpenStudy (anonymous):

oh ok ill fix it then thx

OpenStudy (anonymous):

1 mi = 1609.34 m

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

but i do understand what you did and thanks

OpenStudy (anonymous):

this dimensional analysis stuff does not agree with me at times lol

OpenStudy (anonymous):

ok then make the correction and it'll look like this time = 240000*1609.34/(3*10^8)

OpenStudy (anonymous):

haha kk :)

OpenStudy (anonymous):

n of course multiply the answer by 2

OpenStudy (anonymous):

you could just begin with 480000 right?

OpenStudy (anonymous):

yea u could just because its doesnt throw off the working

OpenStudy (anonymous):

480000 mi * 1609.34m/mi * s/300000000

OpenStudy (anonymous):

300000000 m oops

OpenStudy (anonymous):

if this is home work then you shouldnt multiply by 2 first actually cause yes you'll get the rigt answer but when u write ur statements u'll have to explain y u muliplied by 2 n it makes more sense to do it after

OpenStudy (anonymous):

i still get the wrong answer i think >.< i get 2.57 s

OpenStudy (anonymous):

yea that is the correct answer

OpenStudy (anonymous):

(4800000 * 1609.34)/300000000 = 2.57s

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

k thanks a ton

OpenStudy (anonymous):

sure thing :)